OpenStudy (anonymous):

Please Help Me!! Will Fan! So i kind of understand how to find the trigonometric functions but i really need help on writing the domain (sin theta) (sec theta)= (tan theta)

OpenStudy (anonymous):

the problem is labeled as number 8, i dont know how to write the domain. this is not my work

jimthompson5910 (jim_thompson5910):

what is sec(theta) in terms of sine or cosine?

OpenStudy (anonymous):

1/cos theta

jimthompson5910 (jim_thompson5910):

because we have a fraction with a variable term in the denominator, this means there is a possibility that the denominator could be 0

jimthompson5910 (jim_thompson5910):

when is the denominator zero?

OpenStudy (anonymous):

never!

OpenStudy (anonymous):

or else it'll be undefined, right?

OpenStudy (anonymous):

i really dont understand how to find the domain of validity

jimthompson5910 (jim_thompson5910):

you have the right idea the denominator *should* not be zero to make sure you're not dividing by zero (since it's undefined) but there is a possibility of a variable expression being equal to zero

jimthompson5910 (jim_thompson5910):

for example 1/(x-6) is undefined when x = 6 because this value makes the denominator zero

OpenStudy (anonymous):

oh i see

jimthompson5910 (jim_thompson5910):

if we restrict that x cannot equal 6, then we avoid this division by zero error

jimthompson5910 (jim_thompson5910):

similarly, cos(theta) could be equal to zero at certain values of theta

jimthompson5910 (jim_thompson5910):

when is cos(theta) = 0 true?

OpenStudy (anonymous):

im not too sure, does it equal zero at 1?

jimthompson5910 (jim_thompson5910):

do you have a unit circle?

OpenStudy (anonymous):

oh, is it pi/2?

OpenStudy (anonymous):

because that's the first radian on the cosine graph?

jimthompson5910 (jim_thompson5910):

good, cos(pi/2) = 0

jimthompson5910 (jim_thompson5910):

what other values do you find?

OpenStudy (anonymous):

pi, 3pi/2, and 2pi

jimthompson5910 (jim_thompson5910):

3pi/2, yes 2pi, no

jimthompson5910 (jim_thompson5910):

cos(2pi) = 1

OpenStudy (anonymous):

5pi/2?

jimthompson5910 (jim_thompson5910):

so what's a general rule to capture all of the solutions of cos(theta) = 0

OpenStudy (anonymous):

all odd multiples of pi/ 2

jimthompson5910 (jim_thompson5910):

that's one way, what's another?

OpenStudy (anonymous):

all reals except odd multiples of pi/2?

jimthompson5910 (jim_thompson5910):

well another way is to start with pi/2 then add on multiples of pi (so pi/2 + pi = 3pi/2 or pi/2 + 2pi = 5pi/2, etc etc)

jimthompson5910 (jim_thompson5910):

so one complete solution set is \[\Large \theta = \frac{\pi}{2} + \pi*n\] where n is any integer

OpenStudy (anonymous):

ohh, okay that makes a lot of sense

jimthompson5910 (jim_thompson5910):

going back to the original equation, theta can be any value you want but it cannot be values in the form of \[\Large \theta = \frac{\pi}{2} + \pi*n\] where n is any integer

jimthompson5910 (jim_thompson5910):

this is to make sure you don't divide by zero